Thursday, May 24, 2007
Dental School :
- Zimmer started the pre-cleaning and wipe down in the rooms of the 6th floor of the Clinic Building.
- The duct cleaning continues on the 5th floor.
- The follow up cleaning of the 4th and 5th floors continues.
- Boe Tel pulled network cable on the 6th and 7th floors. They continued making the connections on the 4th and 5th floors.
- Schindler worked on the two elevators in the Clinic Building that experienced outages.
- They continued to work on the three elevators in the penthouse of the Administration Building .
1542 Tulane Ave. :
- Ordes Electrical is running wire for the pump breaker panel in the basement.
- Arc Mechanical is installing the new steam condensate pumps.
- Stratos Elevator is installing the travel cable for Elevator #3.
- The demolition of the old terrazzo floor continues.